OpenStudy (anonymous):

find the vertex and the parabola: f(x)=-x^2-2x+3

OpenStudy (anonymous):

*of the parabola

OpenStudy (amistre64):

the vertex is when x = -b/2a

OpenStudy (amistre64):

--2/2.1 = 2/2 = 1 (1,y)... -1-2+3=0 (1,0) is the vertex

OpenStudy (amistre64):

i think that typoed...

OpenStudy (amistre64):

--2/2.-1 = 2/-2 = -1

OpenStudy (amistre64):

(-1,y)...-1+2+3 = 4 (-1,4)

OpenStudy (mathmagician):

vertex is (-1,4)

myininaya (myininaya):

you could always write it in vertex form f(x)=-(x^2+2x)+3=-(x^2+2x+1)+3+1=-(x+1)^2+4 vertex is (-1,4)